AnimeGhoul8863:

Joselyn is a manager at a sign-painting company. She has two painters, Allen and Brianne. Allen can complete a large project in 16 hours. Brianne can complete the project in 18 hours. Joselyn wants to know how long it will take them to complete the project together. Write an equation and solve for the time it takes Allen and Brianne to complete the project together. Explain each step.

AnimeGhoul8863:

Formula: \[t=\frac{ AB }{ A+B }\]

AnimeGhoul8863:

\[t \frac{ (16)(18) }{ 16+18 }\]

AngeI:

That looks right e.e

AnimeGhoul8863:

\[t=\frac{ (16)(18) }{ 34 }\]

AngeI:

Keep going

AnimeGhoul8863:

\[t=\frac{ 288 }{ 34 }\] now we going to make a mixed fraction bare with meh

AngeI:

Couldnt you just divide

AnimeGhoul8863:

\[8=\frac{ 16 }{ 34 }\]

AnimeGhoul8863:

e.e i tried but it gave me a decimal so i subtracted instead sorry

AnimeGhoul8863:

but is this right

AngeI:

I dont even know anymore, im confused

AnimeGhoul8863:

ok hold on e.e

AnimeGhoul8863:

For this equation we are going to use the formula \[t=\frac{ AB }{ A+B }\] A=Allen, B=Brianne, t=Time We then plug in the hours it takes then to finish the job on there own \[t=\frac{ (16)(18) }{ 16+18 }\] we then add the denominator together \[t=\frac{ (16)(18) }{ 34 }\] we then multiply the numerator together \[t=\frac{ 288 }{ 34}\] we then turn it into a mixed fraction to determine the hours and minutes it would take them to do it together \[8=\frac{ 16 }{ 34 }\]

AnimeGhoul8863:

its better to write it \[8+\frac{ 16 }{ 34 }\]

AngeI:

um

AngeI:

@Shadow

AnimeGhoul8863:

\[\frac{ 16 }{ 34 }=\frac{ 8 }{ 17 }\]

AnimeGhoul8863:

\[8+\frac{ 8 }{ 17 }\]

AnimeGhoul8863:

Now we have to convert the 8/17 to minutes using a proportion \[\frac{ 8 }{ 17 }=\frac{ x }{ 60 }\]

AnimeGhoul8863:

\[8+\frac{ 28 }{ 60 }\] this means if Allen and Brianne work together it would take them both 8 hours and 28 minutes to finish the project

AngeI:

That makes sense I think e.e

AnimeGhoul8863:

tank yew angel e.e